OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -1/2x + 5. y = -2x + 2 y = 2x - 14 y = 1/2x - 8 y = -1/2x - 4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what's the slope of y = -1/2x + 5

OpenStudy (anonymous):

how do i find slope

jimthompson5910 (jim_thompson5910):

hint: y = mx+b is your general slope intercept equation

jimthompson5910 (jim_thompson5910):

for y = mx+b, m is the slope

OpenStudy (anonymous):

got it okay give me a minute to work it out

OpenStudy (anonymous):

1/2 right?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

close, -1/2

jimthompson5910 (jim_thompson5910):

the slope of y = -1/2x + 5 is -1/2 any parallel line will also have a slope of -1/2 (parallel lines have equal slopes)

jimthompson5910 (jim_thompson5910):

this parallel line passes through (4, -6) so (x,y) = (4, -6) x = 4 y = -6

OpenStudy (anonymous):

so it would be d right away right???

jimthompson5910 (jim_thompson5910):

y = mx+b y = (-1/2)*x+b ... plug in the slope -6 = (-1/2)*4+b ... plug in the given point solve for b

jimthompson5910 (jim_thompson5910):

yes because D is the only answer choice with a slope of -1/2 I still recommend solving for b

OpenStudy (anonymous):

b=8 right

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

-6 = (-1/2)*4+b -6 = -2+b -6+2 = b -4 = b b = -4

jimthompson5910 (jim_thompson5910):

Now plug m = -1/2 and b = -4 into y = mx+b y = mx+b y = -1/2x - 4

jimthompson5910 (jim_thompson5910):

so that shows you why D is correct

OpenStudy (anonymous):

okay thanks for clearing that up i see what i did wrong! :)

jimthompson5910 (jim_thompson5910):

you're welcome